# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 3.5. Security Challenges and Solution Framework (RQ4 and RQ5)
Placement Strategy category highlights the need for secure service deployment on volunteering fog nodes in multi-cloud environment. Of the 121 reviewed studies, 6 proposed solutions for these challenges (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)). The solutions involve developing a security model that integrates SSL/TLS protocols for communication and certificate-based access controls to protect deployed services on fog nodes. The strategies include implementing a combination of security groups, VPCs, and infrastructure-level measures to distribute risks and prevent unauthorized access. The reviewed studies recommend adaptive methods for handling security in dynamic and decentralized edge-cloud environments by proposing cross-technology security updates and resource allocation frameworks to avoid conflicts and vulnerabilities.
[…]
3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
Table LABEL:tab:Automation-solutions presents an overview of the automation challenges and solution framework for containerized applications in multi-cloud environment. The figure consists of eight categories, each focusing on different aspects of automation challenges and their solutions in a multi-cloud environment.
@@ 3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
Multi-Cloud Automation category reports on challenges and solutions related to automating processes for combining and administering various cloud computing environments. Out of the 121 reviewed studies, 13 identified these challenges and provided associated solutions (see Table LABEL:tab:Automation-solutions and Automation Challenges-Solutions sheet in (Waseem et al., 2024)). Key challenges involve the integration, provisioning, orchestration, and adaptation of multiple clouds. The solutions focus on utilizing proxies for environment interfaces to ease management across multiple cloud platforms and employing model-driven engineering techniques to manage provisioning and adaptation more efficiently. Dynamic on-demand fog computing was identified as an essential approach for handling multi-cloud environment dynamically. At the level of cloud redundancy and resource management driven by security policies, smart and user-friendly automation techniques, such as MSSF, were recommended. For managing the complexity of multi-cloud orchestration, serverless deployment methodologies were proposed, using models like TOSCA-based orchestration to improve scalability and coordination between different cloud stacks. Recommendations also include deploying application-level resource managers to support multi-cloud operations and adopting secure abstraction models to create uniform interfaces for managing a wide range of cloud products. MUSA Deployer tools were highlighted for their key role in automating deployment and monitoring activities, ensuring overall security compliance, and simplifying the deployment pipeline across federated clouds.
@@
Automation in Deployment and Scaling category reports on challenges and solutions related to automating deployment and scaling processes for containerized applications across multi-cloud platforms. Out of the 121 reviewed studies, 12 identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). Key challenges include intelligent container placement, application runtime switch automation, and service deployment issues. The proposed solutions emphasized intelligent algorithms for container placement, using machine learning for multi-cloud workload balancing. Runtime switch techniques are also being automated with methods such as service drivers, which are designed to be lightweight and enable seamless application migration and replication. DevOps methodologies were frequently mentioned in the literature in relation to managing deployment and scaling through continuous development and error reduction. Security during deployment is ensured through automated security measures integrated into the pipeline, aiming for a secure setup with reduced vulnerabilities. The automation strategies include elastic resource provisioning using predictive analytics to handle fluctuating workloads in real time. Optimization models were involved in solutions for virtual function placement and container deployment to enhance efficiency. This emphasis on automating and optimizing deployment strategies led to the proposal of comprehensive multi-cloud orchestration tools that facilitate the deployment of containers and VMs, aimed at improving agility and scalability within diverse cloud environments.
@@
Table 12. Identified automation challenge and solution for containerized applications in multi-cloud environment
@@
(Sabbioni et al., 2020) Integrating multi-cloud environment Implementing Environment Proxies
(Ferry et al., 2014) Automating Multi-Cloud Provisioning Adopting Model-Driven Engineering
(Sami and Mourad, 2020) Creating On-Demand Fog in Multi-Cloud Enabling On-Demand Fog Computing
(Libardi et al., 2014) Enhancing User-Friendly Automation Using Multi-Cloud Storage Selection
(Wurster et al., 2018) Orchestrating Multi-Stack Environments Deploying with TOSCA-based Approach
(Rios et al., 2017) Securing Multi-Cloud Application Creation Applying MUSA Framework
(Dimitrijevic et al., 2019) Unifying Resource Abstraction Managing with Resource Managers
(Casola et al., 2017) Integrating Multi-Cloud Deployments Automating MUSA Deployment
(Tricomi et al., 2017) Coordinating Federated Cloud Deployments Orchestrating Subsystem Deployment
(Sitaram et al., 2018) Setting Up Hybrid multi-cloud environment Standardizing Interoperability
(Quint and Kratzke, 2018) Ensuring Multi-Cloud Interoperability Enabling Multi-Cloud Support
(Challita et al., 2017) Managing Multi-Cloud Brokerage Systems Managing Cloud Services
(Kritikos et al., 2019) Implementing Multi-Cloud Orchestration Migrating Manually to Cloud
(Bou Ghantous and Gill, 2021) Establishing DevOps in Multi-Cloud Automating Agile Development
(da Silva et al., 2023) Addressing Cost and Migration Challenges Managing Complexity
(Pandi et al., 2023) Overcoming Multi-Cloud Automation Barriers Integrating Jenkins and Kubernetes
(Ramesh et al., 2023) Improving Multi-Cloud Cost Efficiency Optimizing Deployment Costs
(Petrovski and Gusev, 2024) Simplifying Multi-Cloud Deployment Complexity Adopting AWS Deployment Strategies
(Kataru et al., 2023) Maintaining Multi-Cloud Applications Automating Resource Allocation
@@
(Alonso et al., 2019) Automating Multi-Cloud Deployment Monitoring Adopting Multi-Cloud Strategy
(Deng et al., 2022) Optimizing Container Placement Implementing Intelligent Placement
(Zhang et al., 2021) Automating Runtime Application Switching Deploying Service Driver Model
(Pham and Pham, 2015) Enabling Fine-Grained Component Automation Replicating and Migrating Services
(Barna et al., 2017) Streamlining DevOps Automation Developing with DevOps Approach
(Gupta et al., 2021) Automating Service Deployment Using Automated Scripts
(Sheridan et al., 2017) Securing Deployment Time Securing with Automation Tools
(Mittal and Risco-Martín, 2017) Automating Distributed Simulations Adopting DevOps Methodologies
(Al-shammari and Alwan, 2018) Improving Software Deployment Speed Optimizing Container Images
(Lorenzo et al., 2019) Automating Container Startup Profiling Container Execution
(Pandey et al., 2022) Dynamically Scaling Resources Automating Scaling with CSP Capability
(Pahl, 2015) Scaling Container Deployment Scaling with Kubernetes
(Saif et al., 2023) Provisioning Elastic Resources Provisioning Resources Elastically
(Bhamare et al., 2015) Automating VF Placement Optimizing VF Placement
(Guerrero et al., 2018a) Managing Container and VM Deployment Orchestrating Multi-Cloud Deployments
@@
(Ramos et al., 2021) Handling Resource Overbooking Detecting Overbooking with ML
(Cui et al., 2016) Managing Virtual Clusters Managing with ClaaS Model
(Altran et al., 2022) Scheduling Containerized Applications Scheduling with Kubernetes
(Shi et al., 2020) Selecting Optimal VM Types Using GA-based Algorithms
(Baresi et al., 2016) Automating Resource Commissioning Adopting MicroCloud Architecture
(Dreibholz et al., 2019) Supporting Cloud Federation Automation Optimizing with Middleware Platform
(Trihinas et al., 2015) Automating Elastic Resource Detection Automating with JCatascopia
(Buyya and Barreto, 2015) Automating Resource Acquisition Managing Resources with Aneka
(Mfula and Nurminen, 2018) Enforcing SLA Compliance Testing for Compliance
@@
(Almeida et al., 2014) Safeguarding Data Integrity in Migration Maintaining Integrity Autonomously
(Slominski et al., 2015) Migrating Legacy Web Applications Reusing Architectural Patterns
(Kiss, 2018) Porting Applications to Cloud Using CloudSME Platform
@@
(Verdugo et al., 2017) Automating User-Oriented Testing Virtualizing with Lightweight Containers
(Jha et al., 2019) Streamlining Benchmarking Orchestrating Docker Benchmarking
(Naik, 2016a) Enhancing Automation Dependability Enhancing Dependability with CI
@@
(Gattobigio et al., 2022) Standardizing Automation Processes Integrating Service Mesh
(Lazuka et al., 2022) Optimizing Cloud Provider Selection Selecting Optimal Providers
(Ouyang et al., 2022) Unifying Microservices Deployment Adopting Service Chain Models
(Harwalkar et al., 2019a) Addressing Cloud Interoperability Managing with Multi-Cloud Harmony
(Pellegrini et al., 2017) Standardizing Cloud Applications Standardizing Interfaces
(Pai and Kumar, 2021) Automating Package Creation and Scaling Automating Package Frameworks
@@
(Bolivar et al., 2018) Selecting 5G Ecosystem Components Utilizing Open-Source Components
(Rahman and Lama, 2019) Managing End-to-End Tail Latency Modeling Performance Efficiently
(Serhiienko and Spillner, 2018) Handling Distributed Application Complexity Managing Distributed Applications
@@
(Alaluna et al., 2017) Addressing Network Complexity Managing Network Complexity
(Wahab et al., 2016) Improving Social Network Service Discovery Automating Service Discovery
(Pham et al., 2015) Administering Complex Applications Building Autonomic Systems
@@
Resource Management Automation category reports on challenges and solutions related to automating the management of resources in multi-cloud environment. Out of the 121 studies reviewed, 9 specifically addressed these challenges and provided appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). These challenges include resource overbooking management, resource management for containerized applications, and resource commissioning and decommissioning. The proposed solutions include machine learning-based resource overbooking detection for efficiently managing service containers across multiple cloud platforms. To manage containerized applications, custom Kubernetes schedulers, such as label-affinity schedulers, were proposed to optimize orchestration and allocation. Additionally, a GA-based algorithm was suggested to reduce the search space of VM types in data centers, simplifying the resource allocation problem within multi-cloud environment. For resource commissioning and decommissioning, proposals such as the MicroCloud architecture enabled fine-grained resource allocation and coordinated adaptation workflows. The solutions also addressed cloud federation and inter-platform portability challenges, utilizing middleware platforms for security enhancement, cost management, and performance optimization. The JCatascopia automated modular monitoring framework was introduced to monitor runtime configurations and detect elasticity actions. Finally, SLA compliance testing automation was proposed to manage resource allocation processes and enforce corrective actions, ensuring that service level agreements are met across cloud environments.
@@
Data and Application Migration Automation: category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment. Among the 121 studies reviewed, 3 identified these challenges along with their proposed solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The main challenges include ensuring data integrity during the migration process and enabling the smooth migration of legacy Web applications to cloud services. The proposed solutions involve introducing autonomous management cycles to reduce complexity and ensure effective deployment across multi-cloud environment. Other solutions include reusable architectural patterns to accelerate the modernization of legacy applications in containerized multi-cloud environment. Additionally, the CloudSME Simulation Platform was highlighted for its ability to accelerate the porting of existing applications to cloud infrastructures, with support for more complex tasks such as billing and resource management.
@@
Testing and Benchmarking Automation category highlights the challenges and solutions related to automating the testing and benchmarking processes for containerized applications in multi-cloud environment. Out of the 121 studies reviewed, 3 specifically addressed these challenges and provided relevant solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The main challenges revolve around automating user-oriented testing and benchmarking across various cloud platforms. The solutions range from just-in-time deployment and streamlined multicomponent prototyping to rapid testing, enabled by container-based lightweight virtualization. Additionally, Smart Docker Benchmarking Orchestrators were proposed for automating benchmarking in multi-cloud environment, thereby enhancing the overall efficiency of testing procedures. The application of Computational Intelligence (CI) was also suggested to improve dependability in automated testing for containerized multi-cloud systems.
@@
Standardization and Interoperability Challenges category reports on the challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments. Out of the 121 reviewed studies, 6 specifically identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The main challenges involve standardizing automation processes, selecting cloud providers, and addressing the lack of unified templates for microservices deployment. The proposed solutions include service mesh networks and multi-cloud cluster federation to standardize the automation processes for containerized applications. AI-driven tools for optimal cloud provider selection were also introduced, which help in selecting the most suitable provider and node based on factors such as cost-effectiveness and runtime efficiency in multi-cloud environment. To address the lack of common templates for microservices, a Service Model approach was proposed to automate the deployment of microservices across multiple cloud platforms, thereby reducing operational complexity. For cloud interoperability, an AI-driven platform, such as multi-cloud Harmony, was developed to facilitate seamless data transfer and communication between disparate cloud environments. In terms of cloud application standardization, the solutions emphasized promoting the adoption of global standards and interfaces to improve interoperability between cloud service providers and middleware platforms. Finally, automation frameworks for package creation and scaling, such as those using machine-readable definition files, were proposed to efficiently manage both virtual machines and Docker containers across multi-cloud environment.
@@
Application and Service Management category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment.Out of the 121 reviewed studies, 2 specifically addressed these challenges and proposed appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The key challenges include the selection of 5G ecosystem components and end-to-end tail latency management in microservice architectures. The proposed solutions involve using open-source software components to customize the orchestration of containers and virtual machines in modern multi-cloud deployments. Additionally, performance modeling approaches were suggested to evaluate the performance of container-level and VM-level data in multi-cloud environment, enabling more efficient latency management and optimized overall performance.
@@
Runtime and Service Discovery category organizes the challenges and solutions related to runtime management and services discovery in multi-cloud environment, particularly in managing the complexity of network control and administration for containerized applications. Out of the 121 reviewed studies, 4 specifically tackled these challenges (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)). The proposed solutions include automating network management to address the increasing complexity of distributed applications across multi-cloud topologies. Additionally, automated discovery and selection algorithms were proposed to tag and manage social network services within containerized multi-cloud systems. Autonomic computing systems were introduced to handle the growing complexity of applications by facilitating knowledge transfer, enabling more effective containerization management through automation.
[…]
3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)

### v3

@@ 3.5 Security Challenges and Solution Framework (RQ4 and RQ5)
Placement Strategy category highlights the need for secure service deployment on volunteering fog nodes in multi-cloud environment. Of the 121 reviewed studies, 6 proposed solutions for these challenges (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]). The solutions involve developing a security model that integrates SSL/TLS protocols for communication and certificate-based access controls to protect deployed services on fog nodes. The strategies include implementing a combination of security groups, VPCs, and infrastructure-level measures to distribute risks and prevent unauthorized access. The reviewed studies recommend adaptive methods for handling security in dynamic and decentralized edge-cloud environments by proposing cross-technology security updates and resource allocation frameworks to avoid conflicts and vulnerabilities.
Takeaway 8: Security in containerized multi-cloud environments poses unique challenges due to the combined complexity of container orchestration and distributed cloud platforms. The reviewed studies show that addressing these issues requires specific strategies—such as secure container placement, decentralized identity management, and container-aware compliance frameworks—not typically needed in single-cloud or non-containerized setups. These solutions highlight the distinct security needs that arise only when containerization is applied across multiple clouds.
3.6 Automation Challenges and Solution Framework (RQ4 and RQ5)
Table LABEL:tab:Automation-solutions presents an overview of automation challenges and corresponding solutions for containerized applications in multi-cloud environments. The framework is organized into eight categories, each addressing a distinct aspect of automation—such as orchestration, deployment, resource management, and standardization. These categories capture recurring challenges and the solution strategies proposed across the selected studies.
@@ 3.6 Automation Challenges and Solution Framework (RQ4 and RQ5)
Multi-Cloud Automation category reports on challenges and solutions related to automating processes for combining and administering various cloud computing environments. Out of the 121 reviewed studies, 13 identified these challenges and provided associated solutions (see Table LABEL:tab:Automation-solutions and Automation Challenges-Solutions sheet in [11]). Key challenges involve the integration, provisioning, orchestration, and adaptation of multiple clouds. The solutions focus on utilizing proxies for environment interfaces to ease management across multiple cloud platforms and employing model-driven engineering techniques to manage provisioning and adaptation more efficiently. Dynamic on-demand fog computing was identified as an essential approach for handling multi-cloud environment dynamically. At the level of cloud redundancy and resource management driven by security policies, smart and user-friendly automation techniques, such as MSSF, were recommended. For managing the complexity of multi-cloud orchestration, serverless deployment methodologies were proposed, using models like TOSCA-based orchestration to improve scalability and coordination between different cloud stacks. Recommendations also include deploying application-level resource managers to support multi-cloud operations and adopting secure abstraction models to create uniform interfaces for managing a wide range of cloud products. MUSA Deployer tools were highlighted for their key role in automating deployment and monitoring activities, ensuring overall security compliance, and simplifying the deployment pipeline across federated clouds.
@@
Automation in Deployment and Scaling category reports on challenges and solutions related to automating deployment and scaling processes for containerized applications across multi-cloud platforms. Out of the 121 reviewed studies, 12 identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]). Key challenges include intelligent container placement, application runtime switch automation, and service deployment issues. The proposed solutions emphasized intelligent algorithms for container placement, using machine learning for multi-cloud workload balancing. Runtime switch techniques are also being automated with methods such as service drivers, which are designed to be lightweight and enable seamless application migration and replication. DevOps methodologies were frequently mentioned in the literature in relation to managing deployment and scaling through continuous development and error reduction. Security during deployment is ensured through automated security measures integrated into the pipeline, aiming for a secure setup with reduced vulnerabilities. The automation strategies include elastic resource provisioning using predictive analytics to handle fluctuating workloads in real time. Optimization models were involved in solutions for virtual function placement and container deployment to enhance efficiency. This emphasis on automating and optimizing deployment strategies led to the proposal of comprehensive multi-cloud orchestration tools that facilitate the deployment of containers and VMs, aimed at improving agility and scalability within diverse cloud environments.
Resource Management Automation category reports on challenges and solutions related to automating the management of resources in multi-cloud environment. Out of the 121 studies reviewed, 9 specifically addressed these challenges and provided appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]). These challenges include resource overbooking management, resource management for containerized applications, and resource commissioning and decommissioning.
@@
Table 12: Identified automation challenges and solutions for containerized applications in multi-cloud environment
@@
[65] Integrating multi-cloud environment Implementing Environment Proxies
[89] Automating Multi-Cloud Provisioning Adopting Model-Driven Engineering
[29] Creating On-Demand Fog in Multi-Cloud Enabling On-Demand Fog Computing
[79] Enhancing User-Friendly Automation Using Multi-Cloud Storage Selection
[140] Orchestrating Multi-Stack Environments Deploying with TOSCA-based Approach
[56] Securing Multi-Cloud Application Creation Applying MUSA Framework
[61] Unifying Resource Abstraction Managing with Resource Managers
[127] Integrating Multi-Cloud Deployments Automating MUSA Deployment
[128] Coordinating Federated Cloud Deployments Orchestrating Subsystem Deployment
[81] Setting Up Hybrid multi-cloud environment Standardizing Interoperability
[50] Ensuring Multi-Cloud Interoperability Enabling Multi-Cloud Support
[94] Managing Multi-Cloud Brokerage Systems Managing Cloud Services
[76] Implementing Multi-Cloud Orchestration Migrating Manually to Cloud
[90] Establishing DevOps in Multi-Cloud Automating Agile Development
[115] Addressing Cost and Migration Challenges Managing Complexity
[117] Overcoming Multi-Cloud Automation Barriers Integrating Jenkins and Kubernetes
[92] Improving Multi-Cloud Cost Efficiency Optimizing Deployment Costs
[100] Simplifying Multi-Cloud Deployment Complexity Adopting AWS Deployment Strategies
[134] Maintaining Multi-Cloud Applications Automating Resource Allocation
@@
[105] Automating Multi-Cloud Deployment Monitoring Adopting Multi-Cloud Strategy
[69] Optimizing Container Placement Implementing Intelligent Placement
[101] Automating Runtime Application Switching Deploying Service Driver Model
[80] Enabling Fine-Grained Component Automation Replicating and Migrating Services
[52] Streamlining DevOps Automation Developing with DevOps Approach
[103] Automating Service Deployment Using Automated Scripts
[55] Securing Deployment Time Securing with Automation Tools
[106] Automating Distributed Simulations Adopting DevOps Methodologies
[112] Improving Software Deployment Speed Optimizing Container Images
[113] Automating Container Startup Profiling Container Execution
[67] Dynamically Scaling Resources Automating Scaling with CSP Capability
[60] Scaling Container Deployment Scaling with Kubernetes
[77] Provisioning Elastic Resources Provisioning Resources Elastically
[47] Automating VF Placement Optimizing VF Placement
[99] Managing Container and VM Deployment Orchestrating Multi-Cloud Deployments
@@
[132] Handling Resource Overbooking Detecting Overbooking with ML
[126] Managing Virtual Clusters Managing with ClaaS Model
[114] Scheduling Containerized Applications Scheduling with Kubernetes
[97] Selecting Optimal VM Types Using GA-based Algorithms
[31] Automating Resource Commissioning Adopting MicroCloud Architecture
[122] Supporting Cloud Federation Automation Optimizing with Middleware Platform
[84] Automating Elastic Resource Detection Automating with JCatascopia
[98] Automating Resource Acquisition Managing Resources with Aneka
[42] Enforcing SLA Compliance Testing for Compliance
@@
[95] Safeguarding Data Integrity in Migration Maintaining Integrity Autonomously
[44] Migrating Legacy Web Applications Reusing Architectural Patterns
[63] Porting Applications to Cloud Using CloudSME Platform
@@
[120] Automating User-Oriented Testing Virtualizing with Lightweight Containers
[96] Streamlining Benchmarking Orchestrating Docker Benchmarking
[148] Enhancing Automation Dependability Enhancing Dependability with CI
@@
[34] Standardizing Automation Processes Integrating Service Mesh
[74] Optimizing Cloud Provider Selection Selecting Optimal Providers
[121] Unifying Microservices Deployment Adopting Service Chain Models
[108] Addressing Cloud Interoperability Managing with Multi-Cloud Harmony
[73] Standardizing Cloud Applications Standardizing Interfaces
[59] Automating Package Creation and Scaling Automating Package Frameworks
@@
[40] Selecting 5G Ecosystem Components Utilizing Open-Source Components
[53] Managing End-to-End Tail Latency Modeling Performance Efficiently
[45] Handling Distributed Application Complexity Managing Distributed Applications
@@
[124] Addressing Network Complexity Managing Network Complexity
[72] Improving Social Network Service Discovery Automating Service Discovery
[86] Administering Complex Applications Building Autonomic Systems
@@
The proposed solutions include machine learning-based resource overbooking detection for efficiently managing service containers across multiple cloud platforms. To manage containerized applications, custom Kubernetes schedulers, such as label-affinity schedulers, were proposed to optimize orchestration and allocation. Additionally, a GA-based algorithm was suggested to reduce the search space of VM types in data centers, simplifying the resource allocation problem within multi-cloud environment. For resource commissioning and decommissioning, proposals such as the MicroCloud architecture enabled fine-grained resource allocation and coordinated adaptation workflows. The solutions also addressed cloud federation and inter-platform portability challenges, utilizing middleware platforms for security enhancement, cost management, and performance optimization. The JCatascopia automated modular monitoring framework was introduced to monitor runtime configurations and detect elasticity actions. Finally, SLA compliance testing automation was proposed to manage resource allocation processes and enforce corrective actions, ensuring that service level agreements are met across cloud environments.
@@
Data and Application Migration Automation: category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment. Among the 121 studies reviewed, 3 identified these challenges along with their proposed solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]). The main challenges include ensuring data integrity during the migration process and enabling the smooth migration of legacy Web applications to cloud services. The proposed solutions involve introducing autonomous management cycles to reduce complexity and ensure effective deployment across multi-cloud environment. Other solutions include reusable architectural patterns to accelerate the modernization of legacy applications in containerized multi-cloud environment. Additionally, the CloudSME Simulation Platform was highlighted for its ability to accelerate the porting of existing applications to cloud infrastructures, with support for more complex tasks such as billing and resource management.
@@
Testing and Benchmarking Automation category highlights the challenges and solutions related to automating the testing and benchmarking processes for containerized applications in multi-cloud environment. Out of the 121 studies reviewed, 3 studies specifically addressed these challenges and provided relevant solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]). The main challenges revolve around automating user-oriented testing and benchmarking across various cloud platforms. The solutions range from just-in-time deployment and streamlined multicomponent prototyping to rapid testing, enabled by container-based lightweight virtualization. Additionally, Smart Docker Benchmarking Orchestrators were proposed for automating benchmarking in multi-cloud environment, thereby enhancing the overall efficiency of testing procedures. The application of Computational Intelligence (CI) was also suggested to improve dependability in automated testing for containerized multi-cloud systems.
@@
Standardization and Interoperability Challenges category reports on the challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments. Out of the 121 reviewed studies, 6 specifically identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]). The main challenges involve standardizing automation processes, selecting cloud providers, and addressing the lack of unified templates for microservices deployment. The proposed solutions include service mesh networks and multi-cloud cluster federation to standardize the automation processes for containerized applications. AI-driven tools for optimal cloud provider selection were also introduced, which help in selecting the most suitable provider and node based on factors such as cost-effectiveness and runtime efficiency in multi-cloud environment. To address the lack of common templates for microservices, a Service Model approach was proposed to automate the deployment of microservices across multiple cloud platforms, thereby reducing operational complexity. For cloud interoperability, an AI-driven platform, such as multi-cloud Harmony, was developed to facilitate seamless data transfer and communication between disparate cloud environments. In terms of cloud application standardization, the solutions emphasized promoting the adoption of global standards and interfaces to improve interoperability between cloud service providers and middleware platforms. Finally, automation frameworks for package creation and scaling, such as those using machine-readable definition files, were proposed to efficiently manage both virtual machines and Docker containers across multi-cloud environment.
Application and Service Management category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment.Out of the 121 reviewed studies, 2 specifically
@@
addressed these challenges and proposed appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]). The key challenges include the selection of 5G ecosystem components and end-to-end tail latency management in microservice architectures. The proposed solutions involve using open-source software components to customize the orchestration of containers and virtual machines in modern multi-cloud deployments. Additionally, performance modeling approaches were suggested to evaluate the performance of container-level and VM-level data in multi-cloud environment, enabling more efficient latency management and optimized overall performance.
@@
Runtime and Service Discovery category organizes the challenges and solutions related to runtime management and services discovery in multi-cloud environment, particularly in managing the complexity of network control and administration for containerized applications. Out of the 121 reviewed studies, 4 specifically tackled these challenges (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]). The proposed solutions include automating network management to address the increasing complexity of distributed applications across multi-cloud topologies. Additionally, automated discovery and selection algorithms were proposed to tag and manage social network services within containerized multi-cloud systems. Autonomic computing systems were introduced to handle the growing complexity of applications by facilitating knowledge transfer, enabling more effective containerization management through automation.
Takeaway 9: Automation challenges in containerized multi-cloud environments are not merely generic cloud or container issues; they arise from the complex interplay between container portability and cloud heterogeneity. Studies reveal that managing orchestration, scaling, and resource allocation across diverse cloud platforms requires specialized solutions, such as TOSCA-based orchestration, intelligent placement algorithms, and DevOps-integrated pipelines, highlighting how containerization uniquely amplifies automation demands in multi-cloud settings.
3.7 Deployment Challenges and Solution Framework (RQ4 and RQ5)